# Loss Functions and Metrics in Deep Learning

### v3

@@ Abstract
When training or evaluating deep learning models, two essential parts are picking the proper loss function and deciding on performance metrics. In this paper, we provide a comprehensive overview of the most common loss functions and metrics used across many different types of deep learning tasks, from general tasks such as regression and classification to more specific tasks in Computer Vision and Natural Language Processing. We introduce the formula for each loss and metric, discuss their strengths and limitations, and describe how these methods can be applied to various problems within deep learning. We hope this work serves as a reference for researchers and practitioners in the field, helping them make informed decisions when selecting the most appropriate loss function and performance metrics for their deep learning projects.
@@ 1 Introduction
This paper is structured as follows. Section 2 outlines the difference between loss functions and performance metrics. Section 3 provides an overview of the most widely used losses and metrics in regression and classification. Section 4 delves into the prevalent computer vision tasks, detailing the associated loss functions and metrics. Section 5 concentrates on the use of losses and metrics in NLP. Section 6 presents a discussion of this research and, finally, Section 7 rounds out the paper with a concluding statement.
2 Loss Functions vs. Performance Metrics
@@ 2.1 Properties of loss functions
Regression MSE (3.1.1)
MAE (3.1.1)
@@
MSE (3.1.1)
MAE (3.1.1)
RMSE (3.1.2)
MAPE (3.1.2)
SMAPE (3.1.2)
@@
Binary Classification BCE (3.2.1)
@@
PR Curve (3.2.2)
Multi-Class Classification CCE (3.2.1)
Sarse CCE (3.2.1)
@@
Deep Learning Task Loss Functions Performance Metrics
@@
IoU loss (4.2.1)
@@
AP (4.2.2)
AR (4.2.2)
Semantic Segmentation CCE
IoU loss (4.2.1)
@@
IoU (4.2.1),
Pixel Accuracy (4.3.2),
AP (4.2.2)
BF (4.3.2)
Instance Segmentation CCE (3.2.1)
IoU loss (4.3.1)
@@
AP (4.2.2)
Panoptic Segmentation CCE (3.2.1)
@@
PSNR (4.5.2)
SSIM (4.5.2)
IS (4.5.2),
FID (4.5.2)
@@
Deep Learning Task Loss Functions Performance Metrics
Text Classification CE (3.2.1)
@@
Language Modeling T-CE (5.1.1) Perplexity (5.3.6)
BLEU (5.3.4)
@@
Machine Translation T-CE (5.1.1)
@@
Name Entity Recognition T-CE (5.1.1) Accuracy (5.3.1)
P/R/F1 (5.3.2)
@@
Part-of-Speech Tagging T-CE (5.1.1) Accuracy (5.3.1)
P/R/F1 (5.3.2)
Sentiment Analysis CE (3.2.1)
@@
Text Summarization T-CE (5.1.1)
@@
Question Answering CE (3.2.1)
@@
Language Detection CE (3.2.1)
@@
P/R/F1 (5.3.2)
3 Basic Tasks
@@ 3.2.2 Classification Performance Metrics
    Calculate the TPR and FPR at each threshold value from 0 to 1 (see Sections 3.2.2 and 3.2.2).
@@ 4.1 Image Classification
Image classification is used in a diverse range of sectors, demonstrating its broad applicability. In medical imaging, it is used to improve diagnostic accuracy by automatically detecting and classifying abnormalities such as tumors, fractures, and pathological lesions in various types of scans, including X-rays, MRIs, and CT images [68, 69, 70, 71]. In agriculture, it helps to monitor crop health, predict yields, and detect plant diseases through images captured by drones or satellites [72, 73, 74, 75]. The retail industry uses this technology for automatic product categorization, inventory management, and online shopping enhanced with visual search capabilities [76, 77, 78, 79]. Security and surveillance systems use image classification to analyze footage for unauthorized activities, manage crowds, and identify individuals or objects of interest [80, 81, 82]. In environmental monitoring, it is instrumental in tracking deforestation, monitoring changes in the natural environment, and assessing water quality using satellite imagery [82, 83, 84, 85, 86]. In addition, in the manufacturing sector, image classification automates quality control processes by identifying defects and verifying assembly, ensuring that products meet stringent standards with minimal human intervention [87, 88, 89, 90].
@@ 5.3.2 Precision, Recall, and F1 Score
Since we have introduced these metrics before in sections 3.2.2, 3.2.2, and 3.2.2, we will focus on their use for NLP tasks.

### v4

@@ Abstract
When training or evaluating deep learning models, two essential parts are picking the proper loss function and deciding on performance metrics. In this paper, we provide a comprehensive overview of the most common loss functions and metrics used across many different types of deep learning tasks, from general tasks such as regression and classification to more specific tasks in Computer Vision and Natural Language Processing. We introduce the formula for each loss and metric, discuss their strengths and limitations, and describe how these methods can be applied to various problems within deep learning. This work can serve as a reference for researchers and practitioners in the field, helping them make informed decisions when selecting the most appropriate loss function and performance metrics for their deep learning projects.
@@ 1 Introduction
This paper is structured as follows. Section 1.1 outlines the difference between loss functions and performance metrics. Section 2 provides an overview of the most widely used losses and metrics in regression and classification. Section 3 delves into the prevalent computer vision tasks, detailing the associated loss functions and metrics. Section 4 concentrates on the use of losses and metrics in NLP. Section 6 presents a discussion of this research and, finally, Section 7 rounds out the paper with a concluding statement.
1.1 Loss Functions vs. Performance Metrics
@@ 1.2 Properties of loss functions
Regression Mean Squared Error (MSE) (2.1.1)
Mean Absolute Error (MAE) (2.1.1)
@@
Mean Squared Error (MSE)(2.1.1)
Mean Absolute Error (MAE)(2.1.1)
Root Mean Squared Error (RMSE)(2.1.2)
Mean Absolute Percentage Error (MAPE) (2.1.2)
Symmetric MAPE (SMAPE) (2.1.2)
@@
Binary Classification Binary Cross Entropy (BCE) (2.2.1)
@@
Precision/Recall Curve (2.2.2)
Multi-Class Classification Categorical Cross Entropy (CCE) (2.2.1)
Sparse Categorical Cross Entropy (SCCE)(2.2.1)
@@
Computer Vision Task Loss Functions Performance Metrics
@@
Balanced L1 (3.2.1)
Intersection over Union loss (3.2.1)
@@
Average Precision (3.2.2)
Averate Recall (3.2.2)
Semantic Segmentation Categorical Cross Entropy (3.3.1)
Intersection over Union loss (IoU) (3.2.1)
@@
Intersection over Union (IoU) (3.2.1)
Pixel Accuracy (3.3.2)
Average Precision (AP) (3.2.2)
Boundary F1 Score (3.3.2)
Average Recall (AR) (3.2.2)
Instance Segmentation Categorical Cross Entropy (CCE) (3.3.1)
Intersection over Union loss (IoU) (3.2.1)
@@
Average Recall (AR) (3.2.2)
Panoptic Segmentation Categorical Cross Entropy (CCE) (3.3.1)
@@
Peak Signal to Noise Ratio (PSNR) (3.5.2)
Structural Similarity Index (SSIM) (3.5.2)
Inception Score (IS) (3.5.2),
Fréchet Inception Distance (FID) (3.5.2)
@@
NLP Task Loss Functions Performance Metrics
Text Classification Token-level Cross Entropy (4.1.1)
@@
Language Modeling Token-level Cross Entropy (4.1.1) Perplexity (4.3.6)
Bilingual Evaluation Understudy (BLEU) (4.3.4)
@@
Machine Translation Token-level Cross Entropy (4.1.1)
@@
Name Entity Recognition Token-level Cross Entropy (4.1.1) Accuracy (4.3.1)
Precision/Recall/F1 (4.3.2)
@@
Part-of-Speech Tagging Token-level Cross Entropy (4.1.1) Accuracy (4.3.1)
Precision/Recall/F1 (4.3.2)
Sentiment Analysis Token-level Cross Entropy (4.1.1)
@@
Text Summarization Token-level Cross Entropy (4.1.1)
@@
Question Answering Token-level Cross Entropy (4.1.1)
@@
Language Detection Token-level Cross Entropy (4.1.1)
@@
Precision/Recall/F1 (4.3.2)
Retrieval Augmented Generation (RAG) CE (4.1.1)
Contrastive loss (3.4.1)
Marginal Ranking (4.1.4)
NLL (2.2.1)
KL Divergence (3.5.1)
Answer Semantic Similarity (5.2.1)
Answer correctness (5.2.2)
Answer Relevance (5.2.3)
Context Precision (5.2.4)
Context Recall (5.2.5)
Faithfulness (5.2.6)
Summarization score (5.2.7)
Context Entitities Recall (5.2.8)
Aspect Critique (5.2.9)
Context Relevance (5.2.10)
Answer Faithfulness (5.2.11)
Answer Relevance (5.2.12)
2 General Tasks
@@ 2.2.2 Classification Performance Metrics
    Calculate the TPR and FPR at each threshold value from 0 to 1 (see Sections 2.2.2 and 2.2.2).
@@ 3.1 Image Classification
Image classification is used in a diverse range of sectors, demonstrating its broad applicability. In medical imaging, it is used to improve diagnostic accuracy by automatically detecting and classifying abnormalities such as tumors, fractures, and pathological lesions in various types of scans, including X-rays, MRIs, and CT images [68, 69, 70, 71]. In agriculture, it helps to monitor crop health, predict yields, and detect plant diseases through images captured by drones or satellites [72, 73, 74, 75]. The retail industry uses this technology for automatic product categorization, inventory management, and online shopping enhanced with visual search capabilities [76, 77, 78, 79]. Security and surveillance systems use image classification to analyze footage for unauthorized activities, manage crowds, and identify individuals or objects of interest [80, 81]. In environmental monitoring, it is instrumental in tracking deforestation, monitoring changes in the natural environment, and assessing water quality using satellite imagery [81, 82, 83, 84, 85]. In addition, in the manufacturing sector, image classification automates quality control processes by identifying defects and verifying assembly, ensuring that products meet stringent standards with minimal human intervention [86, 87, 88, 89].
@@ 4.1.3 Cosine Similarity Loss
4.1.4 Marginal Ranking Loss
Marginal Ranking loss [260, 261], also known as RankNet loss or pairwise ranking loss, is a loss function commonly used in NLP for tasks that involve ranking or ordering items based on their relevance to a given query or context. The objective of Marginal Ranking Loss is to ensure that relevant items are ranked higher than irrelevant ones, optimizing the model’s ability to distinguish between more and less relevant text segments.
This loss is used in applications where the goal is to learn a ranking function that assigns higher scores to relevant items compared to irrelevant ones. Given a set of items, the model is trained to correctly rank a relevant item (positive sample) higher than an irrelevant item (negative sample) with respect to a query or context.
Formally, let $f(\mathbf{q},\mathbf{p})$ be the model’s score for a positive pair (query $\mathbf{q}$ and relevant item $\mathbf{p}$) and $f(\mathbf{q},\mathbf{n})$ be the score for a negative pair (query $\mathbf{q}$ and irrelevant item $\mathbf{n}$). The Marginal Ranking Loss is defined as (115):
$L(\mathbf{q},\mathbf{p},\mathbf{n})=\max(0,\text{margin}-f(\mathbf{q},\mathbf{p})+f(\mathbf{q},\mathbf{n})),$ (115)
where the margin is a hyperparameter that defines the desired difference between the scores of the positive and negative pairs. The loss is non-zero only when the score for the positive pair is not sufficiently higher than the score for the negative pair. The model is thus penalized for ranking irrelevant items higher than relevant ones, or for ranking relevant and irrelevant items too closely.
Marginal Ranking loss is commonly used in the following NLP tasks.
  1. 1.
    In information retrieval, documents or passages are ranked based on their relevance to a user’s query. Marginal Ranking loss is used to train models to rank relevant documents higher than irrelevant ones, improving the quality of search results. For example, in a search engine, this loss function helps ensure that documents most relevant to the query are placed at the top of the search results.
  2. 2.
    In question-answering systems, Marginal Ranking loss can be used to rank potential answers or passages based on their relevance to the question. The model is trained to score correct answers higher than incorrect or less relevant ones, ensuring that the most accurate and informative answers are provided.
  3. 3.
    For personalized recommendations, Marginal Ranking loss can be applied to rank items (e.g., articles, products) according to their relevance to a user’s preferences. The model is optimized to recommend items that the user is more likely to find relevant, based on past interactions or contextual information.
  4. 4.
    In tasks where the goal is to identify the most similar or relevant text pairs, such as paraphrase detection, Marginal Ranking loss helps the model to correctly rank paraphrases or similar sentences higher than non-paraphrases or dissimilar sentences.
The margin parameter is critical for the effectiveness of Marginal Ranking Loss. A properly chosen margin ensures that the model differentiates well between relevant and irrelevant items. However, setting the margin too high or too low can lead to suboptimal performance, either by making the loss too easy to minimize or by making it difficult for the model to achieve the desired ranking.
Some limitations of this loss include
  • Marginal Ranking loss does not account for situations where positive and negative samples are nearly indistinguishable in relevance. In such cases, the loss function may not strongly penalize the model for minor ranking errors, which could affect performance in scenarios where fine-grained ranking is important.
  • In large-scale ranking tasks, computing Marginal Ranking Loss for all possible pairs can be computationally expensive. Efficient sampling strategies or approximations may be necessary to make the training process scalable.
  • Marginal Ranking Loss is inherently a pairwise approach, meaning that it compares pairs of items at a time. Although effective, this pairwise nature can lead to complexity when dealing with large datasets or when trying to incorporate context from multiple related queries or items.
In recap, Marginal Ranking Loss is used for training models to rank items according to their relevance. It is effective in tasks such as information retrieval, question-answering, and recommendation systems, where the goal is to ensure that the most relevant items are prioritized.
@@ 4.3.2 Precision, Recall, and F1 Score
Since we have introduced these metrics before in sections 2.2.2, 2.2.2, and 2.2.2, we will focus on their use for NLP tasks.